OpenStudy (anonymous):

Which correctly explains how to solve this problem? A number is multiplied by 2. Eight is added to the product. This sum is then divided by 5. The answer is 8. What was the original number? A. Start with 8. Multiply 8 by 5. Then add 8 and multiply the answer by 2. B. Start with 8. Add 8 and 5. Then subtract 8 and multiply the answer by 2. C. Start with 8. Multiply 8 by 5. Then subtract 8 and divide the answer by 2.

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@inkyvoyd

OpenStudy (anonymous):

@flower2468

OpenStudy (anonymous):

coc please do not give out direct answers but guide me to the correct answers

OpenStudy (anonymous):

@mathmale

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

hi

OpenStudy (misty1212):

A number is multiplied by 2. Eight is added to the product. This sum is then divided by 5. The answer is 8. lets call the number \(N\) for "number"

OpenStudy (misty1212):

how to you write "N times 2" in math?

OpenStudy (anonymous):

N*2

OpenStudy (misty1212):

yes, although most math teachers will write "\(2N\)" but same thing

OpenStudy (anonymous):

oh,okay

OpenStudy (misty1212):

then how do you write \(2N\) plus 8?

OpenStudy (anonymous):

2N+8

OpenStudy (misty1212):

got it

OpenStudy (anonymous):

um so its c?

OpenStudy (anonymous):

I'm so confused

OpenStudy (misty1212):

slow down, we are not there yet, we are still writing the equation forget about a,b,c,d lets just do it

OpenStudy (misty1212):

how do you write \(2N+8\) divided by 5?

OpenStudy (anonymous):

2N+8/5

OpenStudy (misty1212):

ok so now we know that is equal to 8 so the equation is \[\frac{2N+8}{5}=8\]

OpenStudy (misty1212):

now lets solve it which we do by working backwards (the point of this question)

OpenStudy (anonymous):

Okay, Still lost in confussion

OpenStudy (texaschic101):

keep in mind, the SUM is divided by 5

OpenStudy (anonymous):

thxs tht helped

OpenStudy (misty1212):

don't worry about being lost, you are learning this you already answered all the important questions, don't psych yourself out you are not supposed to know the answer instantly

OpenStudy (anonymous):

okay

OpenStudy (misty1212):

so what are the steps to solve \[\frac{2N+8}{5}=8\]? what would the first step be?

OpenStudy (anonymous):

um

OpenStudy (anonymous):

multiply

OpenStudy (misty1212):

as judge judy says "um" is not an answer yes, multiply both sides by 5 i) multiply by 5

OpenStudy (anonymous):

um, 5*5?

OpenStudy (misty1212):

that would give \[2N+8=8\times 5\\ 2N+8=40\]

OpenStudy (misty1212):

once you have \[2N+8=40\] what would the next step be?

OpenStudy (anonymous):

divide

OpenStudy (misty1212):

actually you could divide, you could divide both sides by 2, but that is an extra step, and not what your math teacher wants you to say

OpenStudy (anonymous):

ok

OpenStudy (misty1212):

so lets start again with \[2N+8=40\] and you want to get N by itself, which will take two steps

OpenStudy (anonymous):

ugh

OpenStudy (misty1212):

first step, subtract 8 from both sides, that puts all the numbers on the right

OpenStudy (anonymous):

um...........

OpenStudy (anonymous):

Puzzled* Lost in confussion*

OpenStudy (anonymous):

I'll just guess

OpenStudy (anonymous):

I think it's a

OpenStudy (misty1212):

you have a choice here, you can think "this is impossible, it is too hard for me, i will complain about being confused, hate math, fail" or you can think " i can learn this, it can't be that hard"

OpenStudy (anonymous):

I think it's a

OpenStudy (misty1212):

or you can just guess what the answer is because it is multiple choice

OpenStudy (anonymous):

a

OpenStudy (misty1212):

ok guess A

OpenStudy (anonymous):

is that correct?

OpenStudy (misty1212):

it is not right, but it is certainly a guess

OpenStudy (anonymous):

b then

OpenStudy (anonymous):

@flower2468

OpenStudy (anonymous):

well she didn't help

OpenStudy (anonymous):

Which strategy best explains how to solve this problem? Kent is planting rows of seeds in the community garden. He plants 5 seeds in his first row, 11 seeds in his second row, and 17 seeds in his third row. Using the same pattern, how many seeds will Kent plant in his 7th row? A. Make a table. Write the numbers 1, 2, 3, 4, 5, 6, 7 in the top row of the table (the number of rows). In the first box on the second row write 5, in the second box write 11, and in the third box write 17. Look for the pattern. Fill in the rest of the table using the same pattern.

OpenStudy (flower2468):

@Qwertty123

OpenStudy (anonymous):

please help me flower the tutor i just had did not help me

OpenStudy (anonymous):

@Qwertty123

OpenStudy (anonymous):

B. Guess and test. Guess 38 seeds in the 7th row. Divide 38 by the number of seeds in each row. If the answer is not 7, revise your guess, and test your answer again. C. Translate into a number sentence. (1 × 5) + (2 × 11) + (3 × 17) = Calculate the number of seeds in each row and add them all up.

OpenStudy (anonymous):

SOMEONE HELP! PLEASE?

OpenStudy (anonymous):

ugh!!!!

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

Never mind I will just guess

OpenStudy (anonymous):

okay It a

OpenStudy (mathmale):

Please post (separately) the problem you're working on. Please be careful about taking out your frustration on others when you don't understand something. As I see it, misty1212 worked hard to help you.